OpenStudy (anonymous):

Current interest rate: 3% I already did 2; but i dont know if this info is needed for 3 2) Being a smart financial planner, you want to figure out how many months it will be until your principal is paid down to $10,000.00. Solve for t and show all of your work. Note that t will be negative because the number of months will decrease the principal. 3) Lastly, you decide to keep track of your loan four times a month instead of monthly. Solve for the adjusted interest rate

OpenStudy (perl):

what formula you are using

OpenStudy (anonymous):

1) A(t)=p(1+r/n)^nt A(t)=39,145(1+ .03/12)^12t

OpenStudy (anonymous):

did you solve it?